Any help with the below question will be greatly appreciated. I can't seem to figure it out...
A bacteria culture started with a count of 480 at 8:00 A.M. and after t hours is expected to grow to f(t)=480(3/2)^t. Estimate the number of bacteria in the culture at noon the same day.

Answer by stanbon(75887)

bacteria culture started with a count of 480 at 8:00 A.M. and after t hours is expected to grow to f(t)=480(3/2)^t. Estimate the number of bacteria in the culture at noon the same day.
-----------
From 8AM to noon is 4 hours, so t=4.
f(4)=480(3/2)^4
f(4)=480(5.0625)
f(4)= 2430 bacteria.
Cheers,
Stan H.